OpenStudy (anonymous):

Landon is standing in a hole that is 5.1 ft deep. He throws a rock, and it goes up into the air, out of the hole, and then lands on the ground above. The path of the rock can be modeled by the equation y = -0.005x2 + 0.41x - 5.1, where x is the horizontal distance of the rock, in feet, from Landon and y is the height, in feet, of the rock above the ground. How far horizontally from Landon will the rock land? A. 66.71 ft B. 33.35 ft C. 15.29 ft D. 7.65 ft

OpenStudy (jit4won):

find the derivative ... Equate it with zero

OpenStudy (jit4won):

you'll get x

OpenStudy (anonymous):

i don't know what you mean

OpenStudy (anonymous):

x=15.29007974,66.70992026, y=-5.1 this is what i get

OpenStudy (inkyvoyd):

Find the vertex, bro.

OpenStudy (anonymous):

how do i find the vertex?

OpenStudy (anonymous):

so the answer i got is wrong?

OpenStudy (inkyvoyd):

@jollysailorbold , wolfram alpha doesn't help people do the learning part of learning, only the checking answers (unless it's integration lol)

OpenStudy (inkyvoyd):

*sexy, do you know how to complete the square?

OpenStudy (anonymous):

no

OpenStudy (inkyvoyd):

@sexy1993 , is this algebra, or calculus you are learning right now?

OpenStudy (anonymous):

algebra

OpenStudy (inkyvoyd):

Why don't we do this. Let's start with the basic quadratic equation, and write it in vertex form. Watch please :)

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

\(ax^2+bx+c=0\) -> divide both sides by a \(x^2+\frac{b}{a}x+\frac{c}{a}=0\) rewrite again \(x^2+2(\frac{b}{2a})+\frac{c}{a}=0\) Everything make sense?

OpenStudy (inkyvoyd):

@sexy1993

OpenStudy (anonymous):

i'm sorry but i am lost i don't understand

OpenStudy (inkyvoyd):

Which line doesn't make sense?

OpenStudy (anonymous):

all of it

OpenStudy (inkyvoyd):

Ok. We start out with an quadratic equation. Now, a, b, and c, are called constant variables, because we know them when we are given a specific equation, but we are generalizing this. Does that make sense? If not, which part?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

Now, x is called the independent variable, and y is called the dependent variable. This is because we don't exactly know values of x and y. But, we can pick values of x, and we have an output "Y", or, y depends on x. Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

So, every quadratic equation can be written in the form \(ax^2+bx+c=y\)

OpenStudy (inkyvoyd):

Now, let's rewrite this equation. @lgbasallote , help; I'm lazy.

OpenStudy (lgbasallote):

why was i called again?

OpenStudy (inkyvoyd):

To show @sexy1993 how to write an equation in vertex form so he can solve the problem.

OpenStudy (anonymous):